### v2

@@ 2.3 Mamba
Mamba is a simplified SSM architecture, drawing insights from the H3 architecture (Fu et al, 2023), which serves as the basis of many SSM architectures. Specifically, the H3 (Fu et al, 2023) block integrates two discrete SSMs to explicitly simulate linear attention. The commonly used SSM architectures alternately stack the H3 block (Fu et al, 2023) and the MLP block prevalent in modern neural networks. Unlike them, Mamba integrates these two blocks to construct the simplified Mamba block. As illustrated in Fig. 2, the Mamba block can be viewed from two distinct perspectives. Firstly, it replaces the first multiplicative gate in the linear attention-like or H3 (Fu et al, 2023) block with an activation function. Secondly, it incorporates the SSM transformation into the primary pathway of the MLP block. The overall architecture of Mamba consists of repeated Mamba blocks interleaved with standard normalization layers and residual connections.
@@
Mamba inherits the linear scalability in sequence length of the state space models, and also realizes the modeling ability of Transformers. Mamba exhibits the significant advantages of two primary types of foundation architectures in CV, i.e. CNNs and Transformers, making it a promising foundation architecture for CV. In contrast to Transformers, which rely on explicitly storing the entire context for context-based reasoning, Mamba utilizes a selection mechanism, which selectively propagates or filters information along the sequence based on the current token. This ensures the modeling process is influenced solely by the past and current inputs along the sequence, thereby adhering to the principles of causality. Therefore, the 1D and causal characteristics of this selection mechanism become focal points for researchers applying Mamba to CV.
3 Visual Mamba Backbone Networks
In this section, we review representative visual Mamba backbone networks to clarify the fundamental principles and innovations of applying Mamba to CV. These backbone networks can be categorized into pure Mamba and hybrid Mamba networks. Pure Mamba networks solely rely on the Mamba architecture for processing visual data, whereas hybrid Mamba networks combine the Mamba architecture with other neural network architectures, such as CNNs and attentions, to capitalize on their complementary advantages.
Figure 3: Visual Mamba blocks, including Vision Mamba (Vim) (Zhu et al, 2024a), Visual State Space (VSS) (Liu et al, 2024g), Efficient Visual State Space (EVSS) (Pei et al, 2024), PlainMamba (Yang et al, 2024a), and LocalMamba (Huang et al, 2024e) blocks. The original Mamba (Gu and Dao, 2023) block is presented as a reference for the advancements in the visual Mamba blocks. The scan techniques and their decoupling results are displayed to the left and above the corresponding blocks, respectively
3.1 Pure Mamba
Vim:
Vim (Zhu et al, 2024a) is a Mamba-based architecture directly operating on image patch sequences similar to ViT (Dosovitskiy et al, 2021). The input image is first transformed into flattened 2D patches, which are then vectorized using a linear projection layer and added with position embeddings to retain spatial information. Following ViT (Dosovitskiy et al, 2021) and BERT (Devlin et al, 2019), a class token is appended to the sequences of patch tokens. The overall token sequence is then fed into the Vim encoder, which is structured as a series of identical Vim blocks. As illustrated in Fig. 3, the Vim block is a Mamba block integrating a backward SSM path alongside the forward one.
VMamba:
VMamba (Liu et al, 2024g) identifies the direction-sensitive challenge when applying Mamba to 2D images, due to the 1D causal attribute of its selection mechanism. The selection mechanism makes Mamba unable to assimilate information from the portion of the unscanned data. Therefore, 1D scanning is not optimal for simultaneously capturing the dependency information across different directions of images, leading to restricted receptive fields.
To address the challenge, VMamba introduces a Cross-Scan Module (CSM). The CSM transforms the input image into sequences of patches along both horizontal and vertical axes (cross-scan), to scan the sequences in four directions: from the top-left to the bottom-right, the bottom-right to the top-left, the top-right to the bottom-left, and the bottom-left to the top-right. Then the resulting four sequences are individually processed using selective SSMs. This operation can also be seen as performing bi-directional selective SSMs along 2D axes. Consequently, each pixel integrates information from all other pixels in four different directions. Finally, all sequences are transformed into their original 2D layout to constitute maps and merged to obtain the output map (cross-merge). The whole process of cross-scan, selective SSM, and cross-merge constitutes the 2D-Selective-Scan (SS2D).
VMamba first transforms the input image into 2D patches. Then they are fed into multiple stages of VMamba, which are comprised of stacked Visual State Space (VSS) blocks, and down-sampling operations (Liu et al, 2022a) are inserted between these stages to construct the overall hierarchical architecture. VMamba introduces two kinds of VSS blocks, the vanilla VSS block as shown in Fig. 3 and VSS block. The vanilla VSS block is similar to the Mamba block but replaces the 1D convolution layer with a 2D depth-wise convolution layer and the selective SSM with SS2D plus a layer normalization (LN) layer. The VSS block resembles the typical Transformer blocks.
PlainMamba:
PlainMamba (Yang et al, 2024a) is crafted as a non-hierarchical architecture to fulfill several objectives: (1) a non-hierarchical structure facilitates multi-level feature fusion, enhancing integration across different scales; (2) it supports effective fusion of multi-modal data; (3) its simpler architecture tends to offer better generalization capabilities; (4) it is amenable to optimization for hardware acceleration.
Initially, the input image is transformed into 2D patch tokens, and combined with position embeddings to preserve spatial information. Unlike ViT (Dosovitskiy et al, 2021), no special token, such as the class token, is used. Then these tokens are processed by a series of identical PlainMamba blocks, as shown in Fig. 3. The PlainMamba block is similar to the Mamba block, except that the 2D depth-wise convolution layer is employed to substitute the 1D convolution layer, and more importantly, the selective scanning mechanism of Mamba is adjusted for adapting the 1D operation to 2D images. Firstly, as zigzag scan depicted in Fig. 4, a continuous 2D scanning technique is utilized to ensure the spatial adjacency of tokens and prevent discontinuity. Secondly, a direction-aware updating technique is proposed to explicitly incorporate relative 2D position information into the selective scanning process.
Due to the lack of predefined ordering in visual data and their inherent spatial dimensions, several pure Mamba networks focus on scanning techniques. Mamba-ND (Li et al, 2024d) aims to extend Mamba to multi-dimensional data including images and videos. It treats the 1D Mamba layer as a black box and explores how to unravel and order multi-dimensional data. Results show that a chain of Mamba layers and simple alternating-directional orderings achieve superior performance. FractalMamba (Tang et al, 2024b) adopts a fractal scanning curve, namely the Hilbert curve, as depicted in Fig. 4, for processing 2D image patches. Differently, Mamba® (Wang et al, 2024a) identifies that the feature artifacts observed in vision Transformers are more pronounced in Vim (Zhu et al, 2024a). To alleviate this issue, Mamba® refines Vim by strategically inserting input-independent tokens, referred to as register (Darcet et al, 2024) tokens, into the visual token sequence. ARM (Ren et al, 2024b) demonstrates that autoregressive pre-training is well-suited to the Mamba architecture, significantly enhancing its visual performance and unlocking its scaling potential.
3.2 Hybrid Mamba
In addition to standalone usages, several studies have explored the integration of Mamba with other architectural paradigms to construct hybrid Mamba networks.
LocalMamba:
LocalMamba (Huang et al, 2024e) identifies a significant limitation in the Vim (Zhu et al, 2024a) and VMamba (Liu et al, 2024g) models, the disrupted dependencies among spatially local tokens during a single scanning process. To overcome this issue, as local sampling depicted in Fig. 4, LocalMamba divides the input image into multiple local windows to perform SSM in different directions as VMamba (Liu et al, 2024g), while also maintaining the global SSM operations. Furthermore, LocalMamba implements a spatial and channel attention module before patch merging to enhance the integration of directional features and reduce redundancy. The LocalMamba block is illustrated in Fig. 3. In addition, it adopts a strategy to select the most effective scan directions for each layer, thereby optimizing computational efficiency.
EfficientVMamba:
EfficientVMamba (Pei et al, 2024) introduces the Efficient 2D Scanning (ES2D) technique, which employs atrous sampling of patches on the feature map to reduce computational burdens. The atrous sampling is illustrated in Fig. 4. ES2D is used to extract the global features, while a parallel convolutional branch is used to extract the local features. Both feature types are then individually processed by a Squeeze-and-Excitation (SE) block. Collectively, the ES2D, the convolutional branch, and the SE block constitute the core components of the Efficient Visual State Space (EVSS) block. The output of the EVSS block is the summation of the modulated global and local features. The EVSS block is shown in Fig. 3. The EVSS blocks form the early stages of the EfficientVMamba, whereas the Inverted Residual blocks (Sandler et al, 2018) invertedly form the later stages.
Some backbone networks introduce the operations from the frequency domain. SiMBA (Patro and Agneeswaran, 2024b) aims to address the instability issue of Mamba scaling to large networks on vision datasets. It uses Mamba for sequence modeling and proposes a new channel modeling technique termed Einstein FFT (EinFFT). The EinFFT applies the Fourier Transform and performs Einstein Matrix Multiplication (EMM) in the frequency domain. Specifically, EMM reorganizes the input and the weight matrix along the channel dimensions into blocks, such that each block is a diagonal matrix for efficient computation. After multiplication, a non-linear activation function is applied to modulate the eigenvalues in the Mamba block to ensure stability following the principle that stability is achieved if all eigenvalues of the evolution matrix are negative real numbers (Oppenheim and Verghese, 2010). Vim-F (Zhang et al, 2024b) utilizes the Fourier transform to transfer the feature map into the frequency domain and adds it to the original feature map, thus mitigating the drawbacks of Vim (Zhu et al, 2024a). Since each point in the frequency feature map depends on the entire original feature map, frequency-domain scanning ensures a comprehensive global receptive field. Moreover, the translation invariance characteristic of the Fourier transform helps to mitigate the inductive bias introduced by the scanning strategy. Vim-F further eliminates the need for position embedding and introduces a new patch embedding method tailored for Mamba, employing overlapping convolutions to model the spatial correlation between tokens. In addition, to balance performance with efficiency, MSVMamba (Shi et al, 2024a) modifies the SS2D scanning used in VMamba (Liu et al, 2024g) by downsampling the image in three directions within SS2D to shorten the sequence length, forming a Multi-Scale 2D (MS2D) scanning strategy. Additionally, each MSVMamba block incorporates a Convolutional Feed-Forward Network (ConvFFN) to enhance channel-wise information exchange and local feature extraction.
@@ 3.2 Hybrid Mamba
Figure 4: Scan techniques, categorized into four groups, i.e., scan direction, scan axis, scan continuity, and scan sampling
3.3 Summary of Key Improvements
3.3.1 Tokenization
To process 2D images, they are first converted into sequences of visual tokens via a stem module, typically comprising a convolution layer followed by a linear projection layer. The addition of positional embeddings is optional, as the SSM operation inherently possesses causal properties. The inclusion of a class token is also optional. Some works insert extra tokens to act as register (Darcet et al, 2024) or to perceive discontinuities between rows or columns. Existing approaches process image sequences by treating them as either 1D or 2D structures for the SSM transformations and convolution operations in the Mamba-based blocks. Given the integral role of scanning techniques in these processes, we will systematically categorize and delve into these approaches in detail in the subsequent section.
3.3.2 Scan
@@
The selective scanning mechanism is the key component of Mamba. However, its original design for 1D causal sequences poses challenges when adapting it to non-causal visual data. Considerable research efforts are devoted to addressing these challenges. In the following section, we categorize and discuss these efforts into four main groups: scan direction, scan axis, scan continuity, and scan sampling. This categorization is based on the objectives of the scanning techniques. The scan direction addresses the non-causal characteristics of visual sequences; the scan axis deals with the high dimensionality inherent in visual data; the scan continuity considers the spatial continuity of the patches along the scanning path; the scan sampling divides the full image into sub-images to capture spatial information. The illustration of these four groups is shown in Fig. 4.
Scan Direction:
After unfolding the visual data into 1D sequences, different scan directions can be employed to handle the non-causal nature of these sequences. We first denote the original selective scanning technique, which operates in a single direction, as SD. In (Zhu et al, 2024a), bi-directional sequence modeling is employed to enhance the receptive fields reciprocally. We refer to this approach as BD. In addition to the two common scan directions, supplementary shuffle or reordering strategies can be employed for scanning. We denote the utilization of these strategies with the labels +Shuffle or +Reordering, respectively.
Figure 5: Comparative analysis of the performance and computational complexity across various visual backbone architectures, encompassing Convolution-based methods (Radosavovic et al, 2020; Liu et al, 2022b; Yu and Wang, 2024), Transformer-based methods (Touvron et al, 2021; Liu et al, 2021; Yao et al, 2022; Patro and Agneeswaran, 2023; Yuan et al, 2023), and Mamba-based methods (Zhu et al, 2024a; Yang et al, 2024a; Liu et al, 2024g; Pei et al, 2024; Huang et al, 2024e; Chaudhuri and Bhattacharya, 2024; Shi et al, 2024a; Wang et al, 2024a). The symbol size is proportional to the parameter count of respective model, providing a visual indicator of model scale and complexity. Note that all data are accessed from released academic papers to ensure fairness and credibility
Scan Axis:
@@
Visual data differ from typical sequences by possessing 2D or higher-dimensional spatial information that encapsulates both local and global contexts. Current approaches unfold visual data along various axes to thoroughly integrate this spatial information. For instance, the scanning axes for a 2D image typically include the horizontal, vertical, left diagonal, and right diagonal, designated as H, V, LD, and RD, respectively. For 3D visual data, the axes also involve depth or temporal dimensions, denoted as D.
Scan Continuity:
Alternative techniques are employed to handle the 2D spatial information in visual data. Some approaches, inspired by the ViT (Dosovitskiy et al, 2021), flatten images into sequences. This technique, however, may lead to spatial discontinuities between rows or columns; we refer to this technique as Raster scanning. In contrast, other approaches ensure continuous scanning to maintain spatial continuity among adjacent tokens, a technique we denote as Zigzag. Some approaches utilize fractal scanning curves to better capture the structural information within the image. Specifically, the use of the Hilbert curve, which we refer to as Hilbert, preserves their spatial relationships and significantly improves the token adjacency.
Scan Sampling:
We denote sampling techniques that divide the image into sub-images for scanning as Local Sampling and Atrous Sampling based on the original terminology in their papers. Notably, within these sub-images, different combinations of the former three groups of scanning techniques can be applied, allowing for diverse processing approaches without the necessity for uniformity across all sub-images. The comparison of the obtained sub-images and the original full image, which is denoted as Global Sampling, is illustrated.

### v3

@@ 2.4. Mamba Architecture
Mamba-1 is a simplified SSM architecture, drawing insights from the H3 architecture (Fu et al., 2023), which serves as the basis of many SSM architectures. Specifically, the H3 (Fu et al., 2023) block integrates two discrete SSMs to explicitly simulate linear attention. The commonly used SSM architectures alternately stack the H3 block (Fu et al., 2023) and the MLP block prevalent in modern neural networks. Unlike them, Mamba-1 integrates these two blocks to construct the simplified Mamba-1 block, as illustrated in Fig. 2. The Mamba-1 block can be viewed from two distinct perspectives. Firstly, it replaces the first multiplicative gate in the linear attention-like or H3 (Fu et al., 2023) block with an activation function. Secondly, it incorporates the SSM transformation into the primary pathway of the MLP block. The overall architecture of Mamba-1 consists of repeated Mamba-1 blocks interleaved with standard normalization layers and residual connections.
Figure 2. Mamba-1 block: sequential generation of SSM parameters vs. Mamba-2 block: parallel generation of SSM parameters.
[…]
Different from Mamba-1, which maps from $\boldsymbol{x}$ to $\boldsymbol{y}$ under the SSM point of view and sequentially computes $\mathbf{A},\mathbf{B},\mathbf{C}$ after $\boldsymbol{x}$, Mamba-2 produces $\mathbf{A},\boldsymbol{x},\mathbf{B},\mathbf{C}$ in parallel as the SSD layer is regarded as a mapping from $\mathbf{A},\boldsymbol{x},\mathbf{B},\mathbf{C}$ to $\boldsymbol{y}$. Thus the Mamba-2 block simplifies the Mamba-1 block by removing sequential linear projections; namely, SSM parameters $\mathbf{A},\mathbf{B},\mathbf{C}$ are generated at the beginning of the block, independent of the SSM input $\boldsymbol{x}$, as illustrated in Fig. 2. Additionally, a normalization layer as in NormFormer (Shleifer et al., 2021) is added after the multiplicative gate branch to enhance stability.
@@
Mamba inherits the linear scalability in sequence length of the state space models, and also realizes the modeling ability of Transformers. Mamba exhibits the significant advantages of two primary types of foundation architectures in CV, i.e. CNNs and Transformers, making it a promising foundation architecture for CV. In contrast to Transformers, which rely on explicitly storing the entire context for context-based reasoning, Mamba utilizes a selection mechanism, which selectively propagates or filters information along the sequence based on the current token. This ensures the modeling process is influenced solely by the past and current inputs along the sequence, thereby adhering to the principles of causality. Therefore, adapting the inherent 1D and causal characteristics of this selection mechanism to align with the nature of visual data has become a primary focus for researchers applying Mamba to CV.
3. Visual Mamba Backbone Networks
This section reviews representative visual Mamba backbone networks, categorizing them by key architectural advancements and optimization strategies. A comparative analysis of their performance is also presented.
@@ 3. Visual Mamba Backbone Networks
Figure 3. Scanning techniques, categorized into four groups, i.e., scan direction, scan axis, scan continuity, and scan sampling.
[…]
3.1. Architectural Advancements
The typical architecture of sequence models first transforms 2D images into sequences and then applies the sequence transformations. Following this, the architectural advancements of visual Mamba backbone networks focus on tokenization methods, scanning techniques, and architecture designs.
3.1.1. Tokenization
The 2D images are first converted into sequences of visual tokens via a stem module, typically comprising a convolution layer followed by a linear projection layer, to be processed. Given the inherent causal properties of SSM operations, adding positional embeddings to visual tokens is optional.
Some approaches treat these visual tokens as 1D structures, applying 1D SSM transformations and 1D convolution operations, as demonstrated in Vim (Zhu et al., 2024c). Other approaches preserve the 2D structure of visual tokens and adapt operations in Mamba-based blocks to 2D, as seen in VMamba (Liu et al., 2024e). In addition to visual tokens, other token types can be concatenated or added. Some works insert extra tokens to act as registers (Darcet et al., 2024) or to perceive discontinuities between rows or columns.
Differently, GlobalMamba (Wang et al., 2024p) segments images into multiple frequency bands and spatially downsamples them based on their frequency ranges, with lower resolutions for bands in smaller frequency ranges. These bands are then tokenized into causal sequences using a lightweight CNN. Through this, the 2D images are transformed into sequences of causal tokens containing global information.
3.1.2. Scanning Techniques
@@
The selective scanning mechanism is the key component of Mamba. However, its original design for 1D causal sequences poses challenges when adapting it to non-causal visual data. Significant research efforts have focused on developing scanning techniques to overcome these challenges. To provide a clear understanding, we categorize these scanning techniques into four main groups: scan direction, scan axis, scan continuity, and scan sampling, forming a decoupled taxonomy. This categorization also reveals different objectives of the scanning techniques. The scan direction addresses the non-causal characteristics of visual sequences; the scan axis deals with the high dimensionality inherent in visual data; the scan continuity considers the spatial continuity of the patches along the scanning path; the scan sampling divides the full image into sub-images to capture spatial information. The illustration of these four groups is shown in Fig. 3.
Scan Direction
After unfolding the visual data into sequences, different scan directions can be employed to handle the non-causal nature of these sequences.
Single-directional scanning (SD) represents the original selective scanning technique in Mamba (Gu and Dao, 2024), which processes sequences in a single direction along sequence length. Bi-directional scanning (BD) processes the sequences from the forward and backward directions, thus enhancing the receptive fields reciprocally.
Scan Axis
@@
Visual data differ from typical sequences by possessing 2D or higher-dimensional spatial information that encapsulates both local and global contexts. Current approaches unfold visual tokens along various axes of the original visual data to thoroughly integrate this spatial information. For instance, the scanning axes for a 2D image typically include the horizontal (H), vertical (V), left diagonal (LD), and right diagonal (RD). For 3D visual data, the axes also involve the depth (D) dimension. Additionally, Vim-F (Zhang et al., 2024a) applies the Fourier Transform to convert the image into the frequency domain (F), enabling a global receptive field during the scanning.
Beyond the spatial axes, certain approaches apply scanning along the channel (C) dimension to capture additional channel-specific information.
Scan Continuity
Alternative techniques are employed to handle the 2D spatial information in visual data.
Raster scanning flattens images into sequences following the layout of computer memory. However, this technique can cause spatial discontinuities between adjacent rows or columns, as spatial neighboring tokens in the image may become separated in the sequence. In contrast, Zigzag scanning serializes images following a zigzag path, alternating between left-to-right and right-to-left directions across rows (or columns). This ensures continuous scanning by maintaining spatial continuity among adjacent tokens. Some approaches utilize fractal scanning curves to better capture the structural information within the image. Specifically, Hilbert scanning leverages the Hilbert curve, a space-filling curve that recursively subdivides the image and traverses its regions in a structured manner. This technique effectively preserves the spatial relationships between tokens, significantly enhancing their adjacency within the sequence. Reordering scanning divides the sequence into non-overlapping segments and successively samples instances from each segment in a fixed order, ensuring systematic coverage of the entire sequence. Differently, Shuffle scanning randomizes the token sequence order to enhance positional transformation invariance.
Scan Sampling
In addition to operating on the full image, referred to as Global Sampling, various other sampling techniques are employed for specific purposes.
Local Sampling divides the image into sub-images and performs scanning within each, improving the model’s ability to capture local dependencies. Atrous Sampling samples the image into sub-images with skipping, capturing global dependencies while reducing computational complexity. Multi-scale Sampling uses depthwise convolution to generate multi-scale feature maps for scanning. The shorter sequences of low-resolution feature maps help mitigate the long-range forgetting problem while reducing computational cost. Quadtree Sampling uses a quadtree-based image partitioning strategy that estimates the locality score of each token based on its features. Tokens are then adaptively divided into window quadrants at both coarse and fine granularities, enabling the capture of local dependencies across multiple scales. Channel Sampling divides the image channels into groups to improve computational efficiency. Notably, within these sub-regions, different combinations of the former three groups of scanning techniques can be applied, allowing for diverse processing approaches without the necessity for uniformity across all sub-regions.
@@
3.1.3. Architecture
The core components of the visual Mamba backbone network architecture are various blocks, formed by different combinations of the previously discussed tokenization methods, scanning techniques, and the selective SSM transformation.
These network architectures can consist of pure Mamba blocks that solely rely on the Mamba architecture for processing visual data. Vim (Zhu et al., 2024c) is a pure Mamba network structured as a series of identical Vim blocks, which is a Mamba block integrating a backward SSM path alongside the forward one. VMamba (Liu et al., 2024e) is another pure Mamba network. Its block is similar to the Mamba block but replaces the 1D convolution layer with a 2D depth-wise convolution layer and the selective SSM with 2D-Selective-Scan (SS2D). SS2D unfolds the input patches into sequences along four distinct scanning paths (Cross-Scan), processes each sequence using separate selective SSM transformations in parallel, and then reshapes and combines the resulting sequences to obtain the final output map (Cross-Merge). The PlainMamba (Yang et al., 2024a) block is similar to the VMamba block, with two key differences. Firstly, it employs the continuous zigzag scanning technique to maintain spatial adjacency among tokens and prevent discontinuities. Secondly, a direction-aware updating technique is introduced to explicitly incorporate relative 2D position information into the selective scanning process. Differently, Mamba-ND (Li et al., 2024h) treats the 1D Mamba layer as a black box and explores how to unravel and order multi-dimensional data. Its results show that a chain of Mamba layers and simple alternating-directional orderings achieve superior performance.
Figure 4. Visual Mamba blocks, including Vision Mamba (Vim) (Zhu et al., 2024c), Visual State Space (VSS) (Liu et al., 2024e), MSVMamba (Shi et al., 2024a), PlainMamba (Yang et al., 2024a), and LocalMamba (Huang et al., 2024d) blocks. The original Mamba (Gu and Dao, 2024) block is presented as a reference for the advancements in the visual Mamba blocks. The scanning techniques and their decoupling results are displayed to the left and above the corresponding blocks, respectively.
[…]
Table 1. Architectural Advancements for Visual Mamba Backbone Networks.
Methods Tokenization Scan Strategy Architecture Venue Code
Direction Axis Continuity Sampling Block Structure
Vim (Zhu et al., 2024c) 1D BD H Raster Global Pure Plain ICML24
VMamba (Liu et al., 2024e) 2D BD H/V Raster Global Pure Hierarchical NeurIPS24
Mamba-ND (Li et al., 2024h) 1D BD H/V/D Raster Global Pure Plain ECCV24 Oral
LocalMamba (Huang et al., 2024d) 1D BD H/V Raster Local Hybrid Plain/Hierarchical
EfficientVMamba (Huang et al., 2024d) 2D BD H/V Raster Atrous Hybrid Hierarchical
SiMBA (Patro and Agneeswaran, 2024b) 1D SD H Raster Global Hybrid Plain
PlainMamba (Yang et al., 2024a) 2D BD H/V Zigzag Global Pure Plain BMVC24
MSVMamba (Shi et al., 2024a) 2D BD H/V Raster Multi-scale Hybrid Hierarchical NeurIPS24
FractalMamba (Tang et al., 2024c) 2D BD H/V Hilbert Global Pure Plain
Mamba® (Wang et al., 2024j) 1D + Register BD H Raster Global Pure Plain
Vim-F (Zhang et al., 2024a) 1D BD H + F Raster Global Pure Plain
MambaVision (Hatamizadeh and Kautz, 2024) 1D SD H Raster Global Hybrid Hierarchical
GroupMamba (Shaker et al., 2024) 1D BD H/V Zigzag Channel Hybrid Hierarchical
ShuffleMamba (Huang et al., 2024b) 1D BD H Raster/Shuffle Global Pure Plain
StableMamba (Suleman et al., 2024) 1D BD H Raster Global Hybrid Plain
QuadMamba (Xie et al., 2024c) 2D BD H/V Raster Quadtree Pure Hierarchical NeurIPS24
GlobalMamba (Wang et al., 2024p) F-based Global BD H/V Raster Global Pure Hierarchical
The network architectures can also incorporate hybrid Mamba blocks that combine the Mamba architecture with other neural network architectures, such as CNNs and attentions, to capitalize on their complementary advantages. MSVMamba (Shi et al., 2024a) incorporates a convolutional feed-forward network in each block to enhance channel-wise information exchange and local feature extraction. LocalMamba (Huang et al., 2024d) introduces a spatial and channel attention module before patch merging to enhance the integration of directional features and reduce redundancy. SiMBA (Patro and Agneeswaran, 2024b) uses Mamba for sequence modeling and introduces a new channel modeling technique based on Fourier Transforms.
Fig. 4 illustrates a suite of pure and hybrid visual Mamba blocks, including Vision Mamba (Vim) (Zhu et al., 2024c), Visual State Space (VSS) (Liu et al., 2024e), MSVMamba (Shi et al., 2024a), PlainMamba (Yang et al., 2024a), and LocalMamba (Huang et al., 2024d) blocks. The Mamba (Gu and Dao, 2024) block is also included to facilitate a direct comparison, highlighting the evolutionary design of these blocks in the visual domain. The detailed illustrations of scanning techniques of different blocks and their decoupled results presented in Fig. 4 also validate the logic behind our categorization of scanning techniques.
Table 2. Optimization Strategies for Visual Mamba Backbone Networks.
Methods Strategy Venue Code
Mamba-prune (Zhan et al., 2024) Token Pruning NeurIPS24
VSSD (Shi et al., 2024b) Visual State Space Duality
V2M (Wang et al., 2024m) 2D SSM
Spatial-Mamba (Xiao et al., 2024b) Structure-aware State Fusion
SparX-Mamba (Lou et al., 2024) Cross-layer Connection
HRVMamba (Zhang et al., 2024g) Multi-resolution Parallel Design
ARM (Ren et al., 2024a) Autoregressive Pre-training
MAP (Liu and Yi, 2024) Masked Autoregressive Pre-training
Built on different types of blocks, existing architectures also differ in their use of plain or hierarchical structures. Plain structures typically stack identical blocks, while hierarchical structures introduce down-sampling layers between block stacks to generate hierarchical representations.
Details of tokenization methods, scanning techniques, and architecture designs of visual Mamba backbone networks are outlined in Table. 1.
3.2. Optimization Strategies
In addition to architectural advancements, researchers also explore additional optimization strategies to enhance the efficiency and performance of visual Mamba backbone networks. A summary of these strategies is presented in Table. 2. Mamba-prune (Zhan et al., 2024) introduces token pruning into Mamba-based networks for accelerating the computation. VSSD (Shi et al., 2024b) adapts SSD to a non-causal mode to better handle the non-causality of visual data, while V2M (Wang et al., 2024m) and Spatial-Mamba (Xiao et al., 2024b) extend the SSM to 2D forms. SparX-Mamba (Lou et al., 2024) and HRVMamba (Zhang et al., 2024g) introduce additional designs for cross-layer feature aggregation and dense prediction, respectively. ARM (Ren et al., 2024a) and MAP (Liu and Yi, 2024) focus on improving the model performance via pre-training.